OpenStudy (anonymous):

I need explanation : SET THEORY Q.1 This is Cantor's groundbreaking work that provides the mathematical basis for making sense of statements like "infinity plus ten is infinity," or "infinity plus infinite is infinity." Here we understand "infinity" to mean "aleph null," which is used in set theory to denote the cardinality of the set of natural numbers {1, 2, 3, ...}. 6) Explain why the quantity "infinity - infinity" is not defined.

OpenStudy (anonymous):

Since \[\infty + 4 = \infty\]\[\infty+6=\infty\]Then... \[\infty-\infty=4\]\[\infty-\infty=6\]There is no correct way to know.

OpenStudy (anonymous):

Infinity is not a number so normal number operations cannot be said to apply to it.